Classify the conic section of the equation x2 + 2x + y2 + 6y = 15.
Answer by lwsshak3(11628)

Classify the conic section of the equation x2 + 2x + y2 + 6y = 15.
**
x2 + 2x + y2 + 6y = 15
Complete the square:
(x^2+2x+1)+(y^2+6y+9)=15
(x+1)^2+(y+3)^2=15
This is an equation of a circle with center at (-1,-3) with radius=√15.
Standard form: (x-h)^2+(y-k)^2=r^2, ((h,k) being (x,y) coordinates of the center, r=radius